OpenStudy (anonymous):

Will FAN AND MEDAL Rewrite the expression by using grouping symbols. y • 4 + y • z A. y (4 + z) B. 4(y + z) C. z(y • 4) D. y • (4 + y) • z

OpenStudy (anonymous):

Hey @GabeBae

OpenStudy (gabebae):

Hello

OpenStudy (anonymous):

can u help?

OpenStudy (gabebae):

yup I was just looking at it I would pick A

OpenStudy (gabebae):

There's 2 y's

OpenStudy (anonymous):

ok thank you

OpenStudy (gabebae):

y • 4 + y • z = 4y^2z y (4 + z) = 4y+yz = 4y^2z